OpenStudy (anonymous):

Simplify the following expression. 2x − 6y + 3x2 + 7y − 14x thats 3x to the 2 power

OpenStudy (anonymous):

Like terms baby.

OpenStudy (anonymous):

2x - 6y + 3x^2 + 7y - 14x Add like terms 3x^2 - 12x + y

OpenStudy (anonymous):

idk how to do them when i do it turns out wrong

OpenStudy (anonymous):

Like terms are terms that have the same amount of the respective variables. For instance, terms that have x^2 can be added together: x^2 + -5x^2 = -4x^2. If it's something like this though: -x^2y + x^2, you can't add them. x - 2x = -x 1 + 3 = 4 If there are different amounts of different variables, then you can't add them.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):